Encyclopedia Astronautica
2007.11.08 - ISS On-Orbit Status 11/08/07


In preparation for tomorrow's EVA-5, CDR Whitson and FE-1 Malenchenko, before breakfast, took the standard pre-EVA session with the Russian crew health-monitoring program's medical assessment MO-9/Biochemical Urinalysis.

(MO-9 is conducted every 30 days (and also before and after EVAs) and is one of five nominal Russian medical tests adopted by NASA for U.S. crewmembers for IMG PHS (Integrated Medical Group/Periodic Health Status) evaluation as part of the ""PHS/Without Blood Labs"" exam. The analysis uses the sophisticated in-vitro diagnostic apparatus Urolux developed originally for the Mir program. Afterwards, the data are entered in the MEC (Medical Equipment Computer)'s special IFEP software (In-Flight Examination Program).)

Spending time in the Airlock (A/L), Peggy and Yuri completed A/L EL (Equipment Lock) configuration for their spacewalk tomorrow and tagged up with ground specialists at MCC-H to discuss details. As part of last-minute preparations, the spacewalkers installed the regenerated METOX (Metal Oxide) CO 2 absorber canisters in their EMUs (Extravehicular Mobility Units) and underwent the standard pre-EVA PMC (Private Medical Conference).

Dan Tani configured the DCS-760 digital camera to be taken outside and recharged its battery.

FE-1 Malenchenko performed a software upgrade for the BSPN payload server by deleting old files on the RSS1 laptop and installing the new software, uplinked overnight. Afterwards, the new load was tested with a communications check between RSS1 and BSPN. (In preparation for today's upgrade, on 10/29 Malenchenko had conducted the periodic time synchronization between the RSS1 and the BSPN. Before RSS1/BSPN synchronization, the RSS1 is updated with the exact time as per the station clock (which in turn is synchronized daily from RGS/Russian Ground Site). Experiment control application is a payload file transfer program called ShellForKE.)

Yuri prepared three Russian "Pille-MKS" radiation dosimeters, recorded their dosages and equipped each of the two EMUs (Extravehicular Mobility Unit) with a radiation sensor (A0309/CDR & A0310/FE-2). (A third sensor, A0308, was placed in the SM on the PULT reader for background readings.)

At ~1:20pm, the crew conducted its first weekly tagup with the Lead Flight Director at JSC/MCC-H via S-band/audio.

At ~2:50pm EST, Whitson and Malenchenko are beginning their overnight Campout prebreathe and lockout in the A/L for denitrogenation.

EVA-5 will start tomorrow (Saturday) at ~5:55am EST and last an estimated 6h 35min, nominally ending at 12:30pm. Dan Tani will support the spacewalk as IV (Intravehicular) crewmember, sharing in the initial pre-breathe.

(Main objectives of EVA-5 are:

Stow SSPTS (Station-to-Shuttle Power Transfer System) Cable;
Stow PMA-2/Lab Umbilicals;
Remove Lab CETA (Crew & Equipment Translation Aid) Light;
Temp stow Node-2 Port & Starboard Tray Avionics;
Install Node-2 PDGF (Power & Data Grapple Fixture) Horseshoe Connectors & harness;
R&R RPCM 4B-C (Remove & Replace Remote Power Controller Module) on S0 truss;
Remove ACBM (Active Common Berthing Mechanism) Cover/survey CBM;
Mate S0/Node-1 SM Power Cable;
Configure PMA-1/FGB H-Jumpers;
Retrieve BSP (Baseband Signal Processor);
Relocate N 2 VTE (Vent Tool Extender) bag; and
Install Node-2 handrail.)

The FE-2 powered down the ham radio equipment in SM (Service Module) and FGB at ~12:30pm to prevent RF interference with the EMUs during the spacewalk.

Malenchenko performed the regular daily maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM.

Yuri also serviced the SM's ASU toilet facilities systems/replaceables, performing the periodic replacement of the toilet's urine receptacle (M-P) and filter insert (F-V), plus associated hoses and a sensor, and stowing the old units for disposal.

Peggy conducted the weekly 10-min. CWC audit as part of the continuing WDS (Water Delivery System) assessment of onboard water supplies. (Updated "cue cards" based on the crew's water calldowns are sent up every other week.)

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR), RED resistive exercise device (CDR, FE-1, FE-2), and VELO bike with bungee cord load trainer (FE-1).

Dan then transferred the crew's exercise data file to the MEC for downlink, as well as the daily wristband HRM data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Dan also had another 60 minutes for general crew orientation (station familiarization & acclimatization), as is standard daily rule for the first two weeks after starting station residency.

PMA-2 Robotics Preps: At ~8:15-10:15am EST, the MT (Mobile Transporter) was moved by ground commanding from WS-8 (Worksite 8) to WS-4 and the SSRMS (Space Station Remote Manipulator System) was maneuvered to the Lab PDGF (Power & Data Grapple Fixture) pre-grapple position in preparation for the PMA-2 relocation planned for 11/12. Also, the Cupola and Lab RWSs (Robotics Workstations) were to be powered up for tomorrow's EVA-5; the SSRMS will then walk off to the Lab PDGF and prepare for PMA-2 operations on 11/11. DOUG (Dynamic Orbital Ubiquitous Graphics) notes and procedures for the PMA-2 and Node-2 relocations were uplinked for the crew.

CDRA Powerdown: The Carbon Dioxide Removal Assembly (CDRA) in the Lab was deactivated by the ground during 7:00am-12:00pm EST.

RED Cord Replacement: On 11/1, the crew nominally completed changeout of the Resistive Exercise Device (RED) cords. During review of downlinked imagery specialists noted that the cords had been inadvertently installed backwards. The crew was informed to not extend the cord past the 40 inch mark during exercise to preclude damage to the system. The crew reported that they have used RED several times and have not extended the cord past 28 to 32 inches. Ground teams are analyzing forward plans.

CEO (Crew Earth Observation) photo targets uplinked for today were Lake Eyre, Australia(Lake Eyre is a large, usually dry lakebed in south-central Australia and the lowest part of a larger basin that drains nearly a seventh of the continent. It has a complex cycle of filling and drying, not completely linked to the ENSO (El Niņo-Southern Oscillation Cycle). As we are entering a La Niņa this winter and have no recent photographs, CEO would like some baseline imagery of conditions in and around the lake. This pass was to the NE of the lake in afternoon sun), Lake Nasser, Toshka Lakes, Egypt (these man-made lakes are located in the Egyptian Desert west of Lake Nasser and the Nile River. We have monitored their formation and evolution for the past decade. After expanding rapidly westward for several years of high river flow in the Nile, recent images suggest that they are now in retreat. Requested are context mapping views of these lakes or what is left of them), and Tin Bider Impact Crater (this 70-million year old impact site is located in central Algeria just south of the extensive sand sea known as the Grand Erg Oriental. On this pass in early afternoon sun the target was left of track. Since the crater itself is just 6-km in diameter, the ground requested only context views this time).

ISS Orbit (as of this morning, 8:43am EST (= epoch)):

Mean altitude -- 341.9 km

Apogee height -- 343.6 km

Perigee height -- 340.2 km

Period -- 91.37 min.

Inclination (to Equator) -- 51.64 deg

Eccentricity -- 0.0002593

Solar Beta Angle -- -39.7 deg (magnitude decreasing)

Orbits per 24-hr. day -- 15.76

Mean altitude loss in the last 24 hours -- 115 m

Revolutions since FGB/Zarya launch (Nov. 98) -- 51355

More... - Chronology...

Home - Browse - Contact
© / Conditions for Use